# Testing the Molecular Cloud Paradigm for Ultra-High-Energy Gamma Ray Emission from the Direction of SNR G106.3+2.7

### v1

@@ Appendix A Validating the Planck 353 GHz Dust Opacity Map
Two different surveys are used to assess the molecular clouds in this region: the Dame CO survey (Dame et al., 2001) and the 353 GHz Planck dust opacity map (Ade et al., 2011). We use the Dame CO survey to ensure that there is a molecular cloud at the distance of the Boomerang region. Figure 6 shows the velocity-integrated CO map on top, with a black box indicating the Boomerang region, and the velocity range for a given latitude on bottom [^1] https://lweb.cfa.harvard.edu/rtdc/CO/. From the top image, we can see that there is emission in the region and the bottom shows that there isn’t any CO emission in front, or behind, the Boomerang region.
Figure 6: Top: Velocity-integrated CO map (Dame et al., 2001). The black box indicates the Boomerang region. Bottom: Velocity range over a given longitude (Dame et al., 2001). The black box indicates the Boomerang longitude.
From here, we plot the velocity distribution at the location of the Boomerang to see what velocity range should be integrated over in the CO map. Figure 7 shows that about $-20\leavevmode\nobreak\ \mathrm{km/s}$ to $0\leavevmode\nobreak\ \mathrm{km/s}$ is the ideal range. Kothes et al. (2001) found a peak of CO emission at $-6.4\leavevmode\nobreak\ \mathrm{km/s}$, which corresponds to a distance $800\leavevmode\nobreak\ \mathrm{pc}$. We see a similar peak of emission around approximately $-6\leavevmode\nobreak\ \mathrm{km/s}$, which tells us we are seeing a similar CO emission and can use a distance of $800\leavevmode\nobreak\ \mathrm{pc}$ for the SNR and molecular cloud distances. The right shows the integrated CO map for the same velocity range. It can be seen that there is CO emission coincident with the HAWC VHE emission.

### v2

@@ Appendix A Validating the Planck 353 GHz dust opacity map
Two different surveys were used to assess the molecular clouds in this region: the Dame CO survey (Dame et al., 2001) and the 353 GHz Planck dust opacity map (Ade et al., 2011). We used the Dame CO survey to ensure that there is a molecular cloud at the distance of the Boomerang region. Figure 6 shows the velocity-integrated CO map on top, with a black box indicating the Boomerang region, and the velocity range for a given latitude on bottom [^1] https://lweb.cfa.harvard.edu/rtdc/CO/. From the top image, we can see that there is emission in the region and the bottom shows that there isn’t any CO emission in front, or behind, the Boomerang region.
Figure 6: Dame CO survey for the Boomerang region (Dame et al., 2001). Top: Velocity-integrated CO map. The black box indicates the Boomerang region. Bottom: Velocity range over a given longitude. The black box indicates the Boomerang longitude.
From here, we plotted the velocity distribution at the location of the Boomerang to see what velocity range should be integrated over in the CO map. Figure 7 shows that about $-20\leavevmode\nobreak\ \mathrm{km/s}$ to $0\leavevmode\nobreak\ \mathrm{km/s}$ is the ideal range. Kothes et al. (2001) found a peak of CO emission at $-6.4\leavevmode\nobreak\ \mathrm{km/s}$, which corresponds to a distance $800\leavevmode\nobreak\ \mathrm{pc}$. We see a similar peak of emission around approximately $-6\leavevmode\nobreak\ \mathrm{km/s}$, which tells us we are seeing a similar CO emission and can use a distance of $800\leavevmode\nobreak\ \mathrm{pc}$ for the SNR and molecular cloud distances. The right shows the integrated CO map for the same velocity range. It can be seen that there is CO emission coincident with the HAWC VHE emission.